OpenStudy (jaxon):

How do I determine the slope of X - 2y = -3

OpenStudy (anonymous):

solve for y via \[x-2y=-3\] \[x+3=2y\] \[y=\frac{1}{2}x+\frac{3}{2}\] and now it looks like \[y=mx+b\] a line with slope m and y - intercept b

OpenStudy (anonymous):

since you don't really need b you may be able to just eyeball it

OpenStudy (jaxon):

thanks!

OpenStudy (anonymous):

slope=-a/b hence slope=1/2

OpenStudy (anonymous):

subtract x from both sides --> -2y=-x-3 then divide by -2 on both sides --> y= 1/2x +3/2 by y=mx +b the slope is 1/2